40m
QIL
Cryo_Lab
CTN
SUS_Lab
TCS_Lab
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Thu Mar 10 16:29:32 2016, gautam, Update, endtable upgrade, Inventory check 
|
Thu Mar 10 16:32:45 2016, rana, Update, endtable upgrade, Inventory check
|
Mon Mar 14 22:42:23 2016, gautam, Update, endtable upgrade, Inventory check
|
Mon Mar 28 10:43:18 2016, Steve, Update, endtable upgrade, ETMX 4'x2' optical table pictures  
|
Wed Mar 30 17:38:52 2016, gautam, Update, endtable upgrade, X end table proposed layout  
|
Thu Mar 31 09:38:41 2016, Steve, Update, endtable upgrade, Y end 4x3 existing layout
|
Thu Mar 31 19:49:31 2016, gautam, Update, endtable upgrade, proposed layout v2
|
Mon Apr 4 10:59:12 2016, gautam, Update, endtable upgrade, proposed layout v3    
|
Mon Apr 4 15:04:14 2016, gautam, Update, endtable upgrade, COMPONENT REMOVAL
|
Tue Apr 5 11:42:17 2016, gaericqutam, Update, endtable upgrade, TABLE REMOVAL
|
Wed Apr 6 17:52:21 2016, gautam, Update, endtable upgrade, First contact cleaning commenced
|
Thu Apr 7 12:51:24 2016, gautam, Update, endtable upgrade, Beam height differences
|
Thu Apr 7 14:30:58 2016, Steve, Update, endtable upgrade, ETMX table height    
|
Fri Apr 8 09:13:20 2016, Steve, Update, endtable upgrade, ETMX-T beam height 
|
|
Message ID: 12067
Entry time: Thu Apr 7 14:30:58 2016
In reply to: 12066
Reply to this: 12068
|
Author: |
Steve |
Type: |
Update |
Category: |
endtable upgrade |
Subject: |
ETMX table height |
|
|
Quote: |
Steve has finished installing the enclosure on the new endtable. So Eric and I decided to try and lock the X arm and measure the beam height of the transmitted IR beam relative to the endtable. We initially thought of using POX DC as a the LSC trigger but this did not work as there was no significant change in it when the arm was flashing. Eric then tried misaligning the ITM and using AS110 as a trigger - this worked. We then recompiled the ASS model to take AS110 as an input, and ran the dither alignment. After doing so, I measured the beam height at two points on the new endtable.
Bottom line:
- The beam is roughly level across the table (along the North-South direction, within the precision to which I could place the irides and measure the height). The table has also been levelled pretty well...
- The beam height is ~4.7" across the endtable
So the beam is about 0.7" higher relative to the endtable than we'd like it to be. What do we do about this?
- Is it even possible to raise the table by 0.7" so we can have a level beam everywhere? Are there some constraints related to how the enclosure is attached to the window?
- Are we okay with tolerating a solution where we keep the beam level at 4", and use Y10 and Y11 (see layout in elog 12060) to raise the beam by 0.7", and then have slightly higher posts for the optics downstream of this point?
I've also placed two irides extending the cavity axis on the endtable. These should be helpful in aligning the green to the arm eventually.
|
The new TMC 4' x 3' x4" optical table and enclosure is installed - aligned- leveled.
Atm2, Picture is taken ~42" from the window at 3.75 camera height. The leveled table height is wthin 1/4 at the center of the window.
I think this is close enough to move on with the installation of the optics.
We can raise the loaded table in the future if it is needed.
Atm4, Optical table height to floor 33" at the south west corner
Atm3, Enclosure top cover transmission at 1064 nm, 1mm beam size, power level 157 mW, 0 degree incident angle, T 1.3% Metal shield is required above 100 mW hitting the wall of the enclosure!
Atm5, window to enclosure Kapton seal |
|
|
|
|
|